A drain can empty a swimming pool in 8 days. A second drain can empty the same pool in 4 days. In how much time will the two drains working together empty the pool?: A drain can empty a swimming pool in 8 days. A second drain can empty the same pool in 4 days. In how much time will the two drains working together empty the pool?
Answer by jojo14344(809)

Remember:
D[1] = drain 1, 8 days
D[2] = drain 2, 4days
If work together,
=(D[1]+D[2])/2
=(8+4)/2
=12/2 = 6days ---------------> If 2 work together
Thank you,
Jojo